OpenStudy (anonymous):

a parent increases a childs monthly allowance by 20% each year. if the allowance is 8$ per month now how many years will it take to reach 20$ per month? use the equation 20=8(1.2)^x to solve the problem. round to the nearest thousandth

OpenStudy (anonymous):

8 years

OpenStudy (anonymous):

that was wrong

OpenStudy (anonymous):

it is?

OpenStudy (anonymous):

yes it was its ok though lol im just trying to pas this quiz i can take it multiple tmes

OpenStudy (anonymous):

okay well im sorry bout that though

OpenStudy (anonymous):

its all good

OpenStudy (anonymous):

So given the equation 20 = 8(1.2)^x, you have to solve for x. first step: divide both sides by 8 giving 20/8 = 1.2^x or 2.5 = 1.2^x

OpenStudy (anonymous):

uh... that turned out wrong too. Hang on. the answer should be between 5 and 6 I think

OpenStudy (anonymous):

okay, no its right. Are you familiar with the laws of logs?

OpenStudy (anonymous):

What you are going to have to do is put it into log form giving you